alexmath:

Tickets to a baseball game cost $3 for a child’s ticket and $5 for an adult ticket. At the last game, $521 was collected when 139 people attended the game. Which value could replace x in the table so that a single-variable equation can be written and solved to determine the number of child’s tickets, c, sold?

surjithayer:

3c+5x=521 c+x=139 x=139-c substitute the value of x and find the value of c.

alexmath:

thank you

surjithayer:

yw